OpenStudy (anonymous):

Solve and type in a different form by using the given theorems of logarithms. log10M^k =

OpenStudy (anonymous):

\[\log _{10}M^k\]

OpenStudy (jackellyn):

\[\log _{b}(x ^{p})=(p)\log _{b}(x)\] according to this property \[(k)\log _{10}(M)\]